OpenStudy (anonymous):

solve the linear inequality 4-5/6x ≥ 7

OpenStudy (jhannybean):

\[\large \frac{4-5}{6x} \ge 7\] first off,what is 4-5?

OpenStudy (anonymous):

\[4-\frac{ 5 }{ 6 }x \ge7\]

OpenStudy (anonymous):

that is the problem that I need help to solve

OpenStudy (jhannybean):

Ohhh. okay. Misread it.

OpenStudy (anonymous):

\[4-\frac{ 5 }{ 6x }\ge7=-\frac{ 5 }{ 6x}\ge7-4=\frac{ 5 }{6x }\le-3=5\le-18x =-\frac{ 5 }{ 18gex }\]

OpenStudy (anonymous):

does it make a difference that it is \[\frac{ 5 }{ 6 }x \] and not \[\frac{ 5 }{ 6x }\]

OpenStudy (jhannybean):

\[\large 4-\frac{ 5 }{ 6 }x \ge7\] First we will subtract 4 on both sides. \[\large -4 + 4-\frac{ 5 }{ 6 }x \ge 7-4\] Simplify \[\large -\frac56 x \ge 3\] multiply both sides by the negative. And when you multiply or divide by a negative, you flip the sign. \[\large -(-\frac56 x) \le -3 \]simplify \[\large \frac56 x \le-3\]multiply by the reciprocal \[\large \frac65 \cdot \frac56 x \le -3\cdot \frac65\] simplify \[\large x \le \frac{-18}{5}\]

OpenStudy (anonymous):

thank you

OpenStudy (jhannybean):

No problem, do you understand now?

OpenStudy (anonymous):

yes I do

OpenStudy (jhannybean):

awesome :)